# An Evaluation of Real-time Adaptive Sampling Change Point Detection Algorithm using KCUSUM

### 15 Feb 2024 v1

@@ 1. Introduction
A similar topic to change detection is statistical hypothesis testing. In our work, we leverage established methods for two-sample testing and adjust them for change detection. In statistics, two-sample tests are non-parametric evaluations used to determine whether two datasets come from the same distribution. For the task of online change detection, our approach relies on kernel embedding to calculate the Maximum Mean Discrepancy (MMD) statistic. By using the MMD framework, two datasets can be compared by computing the distance between corresponding observable, induced by a positive definite kernel, which is introduced in Section 3. The distance found with MMD is an unbiased estimator and is not restricted to Euclidean datasets. This framework can be further generalized to hypothesis testing that involves strings, graphs or other structured datatypes (Muandet et al., [n. d.]).
Motivated by previous work using the MMD framework in hypothesis testing, we introduce our Kernel Cumulative Sum (KCUSUM) algorithm (Algorithm V.I). In contrast to CUSUM, KCUSUM does not require knowledge of the sample distribution before and after a change occurs. Instead, KCUSUM relies on samples from the pre-change regime and continuously compares them to incoming data using a kernel chosen by the user. By utilizing this framework, KCUSUM can detect changes to any distribution when the distance from the pre-change distribution exceeds a user-defined threshold. Initial theoretical results concern detection delay and average time to false alarm of the proposed algorithm. These results show the derived upper bound for detection time and the expected time until false alarm when no changes are present. Our analysis builds on existing theory for the CUSUM algorithm (Lorden, 1971, 1970).
@@ 2. Cumulative Sum Algorithm
Let us observe a sequence of random variables $X_{n\geq 1}$ with index t such that for all 1 $\leq i<t$, the random variables $X_{i}$ have the distribution $p_{o}$, and for $i\geq t$ have the distribution $p_{1}$. We assume these variables have values in Euclidean space $\mathbb{R}^{d}$, although the proposed kernel variant is not restricted to these parameters. The index at $t$ is the change point with this convention, which is the detection target when performing analysis with the CUSUM algorithm. The analysis is performed in real time, where CUSUM calculates if a change has occurred by time $n$, on data available up to time $n$.
The Cumulative Sum algorithm, or CUSUM, is a live change point detection method proposed by Page (Page, 1954). To properly introduce CUSUM, we assume our two distributions, $p_{o}$ and $p_{1}$, have the densities $f_{o}$ and $f_{1}$, respectively. The full procedure of CUSUM is shown in ALGO II.I. At step $n$, a data point $X_{n}$ is observed, and the log-likelihood ratio $log\frac{f_{o}(x)}{f_{1}(x)}$ is calculated, with the result added onto the statistic $Z_{n}$. If the result is negative, then $Z_{n}$ is set to zero, effectively restarting the algorithm. If Zn crosses a threshold $h$, then a change is declared at time $n$. Formally, the CUSUM stopping time is:
@@
To better see how the CUSUM algorithm works, consider the behavior of the log likelihood ratio before and after the change point t. Before the change, it has mean $-dKL(p_{o},p_{1})$, where $dKL(p_{o},p_{1})=\mathbb{E}_{P_{o}}log\frac{f_{o}(x)}{f_{1}(x)}$ is the Kullback-Liebler divergence between the two distributions. Since $dKL$ is positive when $p_{o}\neq p_{1}$, the increment term will have a negative average. This negative drift combined with the barrier at zero causes the statistic to remain near zero before any change occurs. After the change, the increment has a positive mean equal to $dKL(po,p1)$ and $Z_{n}$ begins to increase, eventually crossing any positive threshold h with probability one, ending the algorithm. Beyond these heuristic arguments, the CUSUM algorithm can be shown to be optimal in a certain sense, as we review below.
@@
Consider a normal distribution $N(a,b)$ with mean a and variance $b$. Let us assume we are detecting a change in the variance of this distribution, with pre-change parameters of $N(1,1)$ and post-change parameters of $N(1,4)$. A sample sequence $X1,...,Xn$ of length $n=400$ with the change in variance at $t=200$ is generated for this example. The change time is marked by a dashed line. The log likelihood ratio in this case is $log\frac{f_{o}(x)}{f_{1}(x)}=\frac{3}{8}x^{2}-\frac{3}{4}x+log\frac{1}{2}+\frac{3}{8}$. It can be seen that this value has a negative mean before the change occurs, and a positive mean after $t=200$. The threshold for CUSUM, $Z_{n}$, gives a detection time at $T_{CUSUM}=212$ with $h=10$.
In this section, we review the performance characteristics of the CUSUM algorithm. Each possible change time defines a different distribution on the sequences $\{X_{n}\}_{n\geq 1}$. If no change occurs, the variables are independent but identically distributed (i.i.d.) with $X_{i}~{}p_{o}$ for all $i\geq 1$. This distribution is denoted on sequences by $\mathbb{P}_{\infty}$, with the respective expectations $\mathbb{E}_{\infty}$. Generally, a change occurring at $t\geq 1$ means that for all $1\leq i\leq t-1,Xi$ is i.i.d. with $X_{i}~{}p_{o}$, and for $i\geq t$, they are i.i.d. with $X_{i}~{}p_{1}$. Let us denote $\mathbb{P}_{t}$ as the probability distribution on sequences under the assumption of a change at time $t$, where $\mathbb{E}_{t}$ represents the expectation for this distribution. For $n\geq 1$, let $F_{n}$ be the $\theta$-algebra $F_{n}=\theta(X_{1},\ldots,X_{n})$. $F_{n}$ represents the information contained in observations up to and including time $n$. Formally, an online change detector can be represented as a stopping time with respect to the filtration $\{F_{n}\}_{n\geq 1}$, that we denote as $T$, which can be interpreted as the estimate for the change point.
When running a change detector on a particular sequence, two types of errors are possible. There may be a false alarm, which means the change is detected too early, or there may be a delay, meaning the change was detected late. We can formalize the levels of false alarm and delay by using the metrics of average run length to false alarm (ARL2FA) and worst case average detection delay (ESADD). These are standard metrics for evaluating change detectors (Lorden, 1971; Moustakides et al., 2009; Poor and Hadjiliadis, 2008).
For a change detector T, the average time to false alarm is:
@@
ARL2FA is the average detection time given a sequence of observations without change. We can measure delay using Lorden’s criterion (Lorden, 1971). If $T$ is a change detector and there is a change at time $t\geq 1$, then the expected delay given the history of the observations up to time $t-1$ is the random variable $\mathbb{E}_{t}[(T-t)^{+}|F_{t-1}]^{1}$ We can obtain the worst case delay for a change at time $t$ by taking the essential supremum over all sequences of length $t-1$, denoted by esssup $\mathbb{E}_{t}[(T-t)+|F_{t-1}]$. By taking the supremum over all change times $t$, we find the worst case delay:
@@
In particular, the CUSUM algorithm provides an optimal trade- off between the time to false alarm (ARL2FA) and the worst case delay (ESADD). This was first proved in an asymptotic form in (Lorden, 1971), and the result was later proved in full non-asymptotic form in (Moustakides, 1986). A proof of optimality can be found in (Moustakides, 1986; Poor and Hadjiliadis, 2008). Further information on the derivation and properties of the CUSUM may be found in (Basseville and Nikiforov, 1993), or (Poor and Hadjiliadis, 2008).
The close relation between the defined threshold $h$ and the performance levels ARL2FA and ESADD is non-trivial and involves solving numerically intractable integrals (Page, 1954; Moustakides, 1986). However, it is possible to derive some upper and lower bounds that may be useful for applications. For the sake of comparison with the analysis of KCUSUM, let us consider the following quantitative bounds on the performance of the CUSUM.
@@
One interpretation of these equations is that by increasing the user defined threshold $h$, we can increase the time until a false alarm at the cost of a longer detection delay. From the ESADD formula, we can see that the detection delay increases when the distributions get closer. The positive term of the log-likelihood ratio is related to the variance of the CUSUM statistic. In Corollary IV.III we shall obtain the somewhat analogous bounds for the KCUSUM algorithm. In the following section, we discuss the Maximum Mean Discrepancy (MMD) statistic before introducing the kernel variant CUSUM algorithm.
@@ 3. Maximum Mean Discrepancy
In statistics, a two-sample test is used to determine whether two datasets are drawn from the same distribution. One method to solve this problem is to consider the empirical measures within the datasets, and compute the distance between measures using probability metrics. If enough data is available, the empirical distance should approximate the true distance. As this distance becomes larger, we become more confident that the datasets come from distinct distributions. This concept is a basic principle underlying several classical testing methods, such as the Kolmogorov-Darling (Feller, 1948), Cramer-von-Mises (Anderson, 1962) and the Anderson-Darling (Pettitt, 1976) tests.
@@
One interesting behavior to note, is that the linear statistic is the sum of i.i.d. terms, which means the central limit theorem may be used to approximate the distribution, which can inform tuning of the thresholds when MMD is used for hypothesis testing (Gretton et al., 2012). Furthermore, in the online setting, we can study the trajectory of the statistic (that is, as a function of $n$) using the theory of sums of i.i.d. random variables, greatly facilitating analysis. We will exploit this structure in the analysis of KCUSUM in the following section.
@@ 4. Related Work
The topic of online change point detection has its roots in industrial quality control systems (Shewhart and Deming, 1939) and sequential analysis (Wald, 1947). In particular, the CUSUM procedure is closely related to the sequential probability ratio test (SPRT), which is a foundational online hypothesis testing algorithm (Wald, 1947). In contrast to online change detection methods, offline detection procedures must wait until the entire sequence is observed, and all data becomes available to make a decision. Kernel offline change point analysis was explored in (Harchaoui et al., 2008). Each theoretical change time is used to partition the dataset into two groups, consisting of prior observations and later observations, and the two resulting datasets are compared using a kernel based test. This is repeated for each possible change time. If one of the comparisons yields a significant discrepancy between the pre- and post- observation datasets, then a change is declared at the point where the largest difference was observed. This procedure is of interest as relies on distributional knowledge, basing its decisions on comparisons between disjoint sets of observations.
Several approaches to non-parametric online change detection have been proposed in (Brodsky and Darkhovsky, 1993, 2010). As shown in their work, they apply only to the case of detecting changes in mean. Another method of applying a kernel non-parametric test to online change detection is the sliding window approach, as pursued in (Li et al., 2015). In this method, at each time a decision regarding the change is made based on the distance between the most recent fixed-size block of data and a block from the reference distribution, using an MMD distance $d_{k}$. This can be seen as the kernel-based generalization of Shewhart control charts (Shewhart and Deming, 1939).
@@
Besides CUSUM, there are other algorithms with optimality properties, most notably the Shiryaev-Roberts (SR) change detector (Fuh, 2003). While the CUSUM algorithm minimizes the worst case delay (ESADD), the SR detector minimizes a form of average delay. Like CUSUM, the SR test statistic takes a simple recursive form, so it may be possible to extend our algorithmic construction to SR-type detectors as well, but this is outside the scope of this paper.
@@ 5. Kernel Cumulative Sum Algorithm
The kernel variant of the Cumulative Sum algorithm (KCUSUM) blends features from the CUSUM procedure with the proposed MMD framework. Instead of taking each new observation $x_{n}$ to compute the log likelihood ratio, which estimates the KL-divergence $dKL$, we use the new observations and reference samples to calculate an estimate of the MMD distance $d_{k}$.
@@
The full step procedure of the KCUSUM algorithm is listed in Algorithm V.1. We assume that the change to be detected is from a reference distribution $p_{o}$ to an unknown distribution $p_{o}$. At even numbered iterations, the current observation $x_{n}$ is paired with the previous observation $x_{n-1}$. Then, the two points are compared with two reference samples $y_{n},y_{n-1}$ with MMD. The constant $\delta>0$ is subtracted from the result to get $v_{n}$. The variable $v_{n}$ is then added to the threshold statistic $Z_{n-1}$ to get the next value of $Z_{n}$. If the new value $Z_{n}$ would be negative, the algorithm resets to 0, effectively restarting the procedure. At odd numbered iterations, the $Z_{n}$ is unchanged. Note that as a consequence, KCUSUM can only raise an alarm at even numbered iterations.
The motivation behind subtracting a positive amount $\delta$ at each step of Algorithm IV.I is to guarantee that the increments of $v_{n}$ have a negative drift under the pre-change regime and positive drift in the post-change regime. This is a consistency property that enables us to formulate non-trivial upper and lower bounds on performance. Using the definitions in Algorithm IV.1, it is evident that before a change, $\mathbb{E}[v_{n}]=-\delta<0$ and after a change, $\mathbb{E}[v_{n}]=d^{2}k(p_{o},p_{1})-\delta$. Theoretically, the algorithm can detect a change to any distribution $p_{1}$ that is at least distance $\surd\delta$ away from the reference distribution $p_{o}$.
@@
Let us consider detecting a change in variance of normally distributed variables, as in Example II.1. We have again the incoming stream of observations and our reference sample, with a change at $t=200$ shown as the red dashed line. The KCUSUM algorithm is based on the linear statistic $P_{L}$, so at each time $n$, the MMD estimate $k(x_{n-1},x_{n})+k(y_{n-1},y_{n})-k(x_{n},y_{n-1})-k(x_{n-1},y_{n})$ is computed. As in the CUSUM, a simple threshold $h$ is used to decide if a change has occurred. For this particular realization of the variables, a threshold of $h=5$ results in detection at $T_{KCUSUM}=225$. The value of $\delta$ was $\delta=\frac{1}{40}$.
@@
Let $T_{KCUSUM}$ be the change detector corresponding to the described KCUSUM procedure (Algorithm IV.1). Let $p_{o}$ be the pre-change, or reference distribution. Then by using a threshold $\delta>0$, the time to false alarm is at least:
@@
If $p_{1}$ is a distribution with $d_{k}(p_{o},p_{1})>\surd\delta$, then the worst case detection delay is at most:
@@
Combining Equations (14) and (16) yields the claim (11). Now we consider bounding the worst case delay. If the sequence $\{x_{i}\}_{i\geq 1}$ has a change at an odd valued time, say $t=2m-1$ for $m\geq 1$, then the sequence $\{Z_{n}\}_{n\geq 1}$ has a change at $t=m$. Explicitly:
(17) $Z_{i}~{}\begin{cases}p_{o}\times p_{o}\times p_{o}\times p_{o}&\text{for $1\leq i<m$}\\ p_{1}\times p_{1}\times p_{o}\times p_{o}&\text{for $m\leq i$}\\ \end{cases}$
@@
The situation is slightly more complex if the change occurs a time t that is even, say $t=2m$ for some $m\geq 1$. In this case, the grouped sequence $\{Z_{n}\}_{n\geq 1}$ does not experience an abrupt change, but instead there are three relevant distributions. Specifically:
(19) $Z_{i}~{}\begin{cases}p_{o}\times p_{o}\times p_{o}\times p_{o}&\text{for $1\leq i<m$}\\ p_{o}\times p_{1}\times p_{o}\times p_{o}&\text{for $i=m$}\\ p_{1}\times p_{1}\times p_{o}\times p_{o}&\text{for $m<i$}\end{cases}$
@@
Combining Equation (21) with the definition of worst case delay shown in Equation (2) yields the inequality shown in Equation (12). We can combine Theorem IV.2 with certain facts about random walks (Lemma A.2 and Proposition A.1) to get more specific bounds on the delays and false alarms, as shown in the following Corollary. In this corollary, we assume that the kernel $k$ is bounded by a constant $||k||_{\infty}$, and also assume that $\delta$ is bounded by $\delta<2||k||_{\infty}$. This is a necessary assumption when the kernel is bounded, since if $\delta\geq 2||k||_{\infty}$, then $g_{\delta}(z)\leq 0$ for all $z$, and it will not be possible to detect any changes.
@@
Let the assumptions of Theorem IV.2 hold. Further assume that the kernel $k$ is bounded by a constant $||k||_{\infty}$ and let $\delta<2||k||_{\infty}$. Then the time to false alarm satisfies:
@@
If $p_{1}$ is a distribution with $d_{k}(p_{0},p_{1})>\surd\delta$, then the worst case detection delay has a maximum of:
(23) $ESADD_{KCUSUM}\leq\frac{2h}{d_{k}(p_{o},p_{1})^{2}-\delta}+\frac{8||k||^{2}_{\infty}}{(d_{k}(p_{o},p_{1})^{2}-\delta)^{2}}$
@@
To upper-bound the last term in this equation, we will apply Lemma A.2. Note that under our assumption that the kernel is bounded, the moment generating function $M(r)=\mathbb{E}_{\infty}[e^{rg_{/delta}(Z_{1})}]$ is guaranteed to be well defined for all $r\geq 0$. Therefore:
@@
where $r$ is any number satisfying $r>0$ and $M(r)\leq 1$. To identify such an $r$, start with a second order expansion of $M$:

### 05 Apr 2024 v2

@@ 1. Introduction
A similar topic to change detection is statistical hypothesis testing. In our work, we leverage established methods for two-sample testing and adjust them for change detection. Generally, two-sample tests are non-parametric evaluations used to determine whether samples are drawn from the same distribution. For the task of online change detection, our approach relies on kernel embedding to calculate the Maximum Mean Discrepancy (MMD) statistic. By using the MMD framework, two datasets can be compared by computing the distance between corresponding observables, calculated by a positive definite kernel, which is introduced in Section 3. The distance found with MMD is an unbiased estimator and is not restricted to Euclidean datasets. This framework can be further generalized to hypothesis testing that involves strings, graphs or other structured datatypes (Muandet et al., [n. d.]).
Since previous work has demonstrated the MMD framework in hypothesis testing, we introduce our Kernel Cumulative Sum (KCUSUM) algorithm (Algorithm V.I). In contrast to CUSUM, KCUSUM does not require knowledge of the sample distribution before and after a change occurs. Instead, KCUSUM relies on samples from the pre-change regime and continuously compares them to incoming data using a kernel chosen by the user. By utilizing this framework, KCUSUM can detect changes to any distribution when the distance from the pre-change distribution exceeds a user-defined threshold. Initial theoretical results concern detection delay and average time to false alarm of the proposed algorithm. These results show the derived upper bound for detection time and the expected time until false alarm when no changes are present. The proceeding analysis of CUSUM relies on prior work for change detection (Lorden, 1970).
@@ 2. Cumulative Sum Algorithm
Let us observe a sequence of random variables $X_{n\geq 1}$ with index t such that for all 1 $\leq i<t$, the random variables $X_{i}$ have the distribution $p_{o}$, and for $i\geq t$ have the distribution $p_{1}$. We assume these variables have values in Euclidean space $\mathbb{R}^{d}$, although the proposed kernel variant is not restricted to these parameters. The index at $t$ is the change point with this convention, which is the detection target when performing analysis with the CUSUM algorithm. The analysis is performed in real time, where CUSUM detects if a change has occurred by time $n$, on observables for a sequence of length $n$.
CUSUM (Cumulative Sum algorithm) is a live change point detection method proposed by Page (Page, 1954). To show how the algorithm detects changes, let us assume two distributions, $p_{o}$ and $p_{1}$, which have the densities $f_{o}$ and $f_{1}$. The full procedure of CUSUM is shown in ALGO II.I. For each step $n$, data ($X_{n}$) is observed, and a log-likelihood ratio $log\frac{f_{o}(x)}{f_{1}(x)}$ is calculated, with each result added to a threshold statistic $Z_{n}$. If a negative value is calculated, $Z_{n}$ becomes 0, which restarts the running calculation. If $Z_{n}$ reaches threshold $h$, then a change has been detected at time $n$. The algorithm ends at time $T_{CUSUM}$:
@@
To better see how the algorithm works, consider the calculation of the log likelihood ratio before and after the change point t. Preceding any change in the distribution, it has mean $-dKL(p_{o},p_{1})$, where $dKL(p_{o},p_{1})=\mathbb{E}_{P_{o}}log\frac{f_{o}(x)}{f_{1}(x)}$ is the Kullback-Liebler divergence between the two distributions. If $dKL$ is positive when $p_{o}\neq p_{1}$, the running total will have a negative average. A negative momentum with the barrier at 0 causes the value to remain near 0 before the distribution changes. If the distribution does change, the value has a positive average $dKL(po,p1)$ and $Z_{n}$ will accumulate the statistic, eventually reaching a threshold h with guaranteed probability, ending the algorithm. Additionally, the CUSUM algorithm is optimized for certain tasks, as we review below.
@@
Consider a normal distribution $N(a,b)$ with mean a and variance $b$. Let us assume we are detecting a change in the variance of this distribution, with pre-change parameters of $N(1,1)$ and post-change parameters of $N(1,4)$. A sample sequence $X1,...,Xn$ of length $n=400$ with the change in variance at $t=200$ is generated for this example. We can then calculate the log likelihood, whose value has a negative mean before the change occurs, and a positive mean afterwards $t=200$. The threshold for CUSUM, $Z_{n}$, gives a detection time at $T_{CUSUM}=212$ with $h=10$.
In this section, we review the performance characteristics of the CUSUM algorithm. Every change point defines a unique distribution on observables $\{X_{n}\}_{n\geq 1}$. If the distribution doesn’t change, the sequences are independent but identically distributed (i.i.d.) with $X_{i}~{}p_{o}$ for all $i\geq 1$. This distribution is denoted on sequences by $\mathbb{P}_{\infty}$, with the respective expectations $\mathbb{E}_{\infty}$. Generally, a change occurring at $t\geq 1$ means that for all $1\leq i\leq t-1,Xi$ is i.i.d. with $X_{i}~{}p_{o}$, and for $i\geq t$, they are i.i.d. with $X_{i}~{}p_{1}$. Assume that $\mathbb{P}_{t}$ is the probability distribution on a sequence with a known change at time $t$, where $\mathbb{E}_{t}$ represents the expectation value for the sequence. For $n\geq 1$, let $F_{n}$ be the $\sigma$-algebra $F_{n}=\sigma(X_{1},..,X_{n})$. $F_{n}$ represents any information for our observations up to $n$. Then, the time of change detection $T$ can be shown with by the filtration $\{F_{n}\}_{n\geq 1}$.
When using a change point detector on any given observables, there are two potential errors that can occur. These errors include a false alarm, where the change is detected prematurely, and a delay, where the change is detected after it has already occurred. By utilizing the metrics of average run length to false alarm (ARL2FA) and worst case detection delay (ESADD), we are able to formalize the levels of false alarm and delay.(Moustakides et al., 2009).
Given any change detection algorithm T, the expected ARL2FA is:
@@
ARL2FA is the mean detection time given observables without change. One method to measure this delay is utilizing Lorden’s criterion (Lorden, 1971). Assume that at time $t\geq 1$, a change occurs. The expected delay with regards to length of the observed sequence up to $t-1$ has the value $\mathbb{E}_{t}[(T-t)^{+}|F_{t-1}]^{1}$. We can show the worst case delay for any change at time $t$ by calculating the essential supremum over all sequences of length $t-1$, given by esssup $\mathbb{E}_{t}[(T-t)+|F_{t-1}]$. We can now calculate ESADD:
@@
In particular, the CUSUM algorithm optimizes the relationship between ARL2FA and ESADD. Further derivation of the proof can be found in (Lorden, 1971). The relationship between thresholds $h$ and ARL2FA and ESADD is non-trivial and requires the determination of complex integration (Page, 1954; Moustakides, 1986). Nevertheless, there are practical upper and lower bounds that can be derived if given a sequence of length $n$. To formalize the analysis for future comparison to KCUSUM, let’s determine some quantitative boundaries for the performance of CUSUM.
@@
We can see that by raising the user defined threshold $h$, we increase the time until a false alarm at the cost of a longer detection delay. Accordingly, one observation of the ESADD formula to note is the behavior of detection delays. As two samples’ distributions approach similar parameters, the delay will increase. Later in this work, we present similar bounds for the KCUSUM algorithm. In the following section, we discuss the Maximum Mean Discrepancy (MMD) statistic before introducing the kernel variant CUSUM algorithm.
@@ 3. Maximum Mean Discrepancy
In statistics, a two-sample test is used to determine whether two datasets are drawn from the same distribution. If enough data is available, the we can assume empirical measurements approximate the true distance within the distribution. As this distance becomes grows, we become more confident that the datasets come from unique distributions. Many classical statistical testing methods, such as the Kolmogorov-Darling (Feller, 1948), Cramer-von-Mises (Anderson, 1962) and the Anderson-Darling (Pettitt, 1976) tests, rely on this principle.
@@ 4. Related Work
The topic of live change point detection was introduced as a method to monitor quality management systems (Shewhart and Deming, 1939) and sequential hypothesis testing (Wald, 1947). In contrast to online change detection methods, offline detection procedures must wait until the entire sequence is observed, and all data becomes available to make a decision. One method of offline change point analysis was explored for information processing systems (Harchaoui et al., 2008), where every possible change time is used to partition the samples into two groups. The two samples consist of pre and post-change observations, where they are compared using various kernels. This process is performed across every partition, which can detect discrepancies between the sample parameters. Then formally, a change is detected at the largest discrepancy. This process only relies on the stream of observables and their direct comparison.
Several approaches to live change detection were suggested in previous research topics (Brodsky and Darkhovsky, 1993; Li et al., 2015). One limitation of this work is that it can only find changes in the mean of the distribution. In many cases, we wish to detect changes in variance or mean, which can be implemented by combining the principles of MMD, with generalized control charts (Shewhart and Deming, 1939). This method utilizes a moving window of predetermined size, similar to convolutions.
@@
Many change detection algorithms exist with various optimal behaviors depending on target use cases. While Cumulative Sum aims to reduce ESADD, Shiryaev-Roberts minimizes average delay (Fuh, 2003). Since both algorithms have a recursive form, it may be possible to extend our research to Shiryaev-Roberts change detectors as well, but the topic is still being investigated.
@@ 5. Kernel Cumulative Sum Algorithm
The kernel variant of the Cumulative Sum algorithm (KCUSUM) combines properties of the CUSUM algorithm with the proposed MMD procedure. Instead of taking each new sample $x_{n}$ to compute a log likelihood, which estimates the KL-divergence $d_{KL}$, we use incoming and reference samples to estimate the MMD distance $d_{k}$.
@@
The full step procedure of the KCUSUM algorithm is listed in Algorithm V.1. Let us assume that a change has occurred where one distribution $p_{1}$, changes to $p_{2}$. During even numbered iterations, the current sample $x_{n}$ is paired with the previous sample $x_{n-1}$. Then, the two points are compared with two reference samples $y_{n},y_{n-1}$ via MMD. The constant $\delta>0$ is subtracted from the result to get $v_{n}$. The variable $v_{n}$ is added to the running threshold statistic $Z_{n-1}$ which becomes $Z_{n}$. If the new value $Z_{n}$ is non-positive, the algorithm restarts. During odd numbered iterations, $Z_{n}$ is not adjusted. Therefore, KCUSUM can only detect changes at even numbered sample indexes.
The motivation behind subtracting $\delta$ between each iteration of KCUSUM is to guarantee a negative pressure on the pre-change data and positive pressure in the post-change calculations. This allows one to formalize non-trivial bounds on the algorithm’s behavior. If we observe the KCUSUM procedure, it can be seen that $\mathbb{E}[v_{n}]=-\delta<0$ before a change occurs, and $\mathbb{E}[v_{n}]=d^{2}k(p_{1},p_{2})-\delta$ after the change. Therefore, KCUSUM can find changes on any distribution $p_{2}$ that is a minimum $\surd\delta$ away from the reference $p_{1}$.
@@
Let us consider detecting a change in the deviation of a normal distribution, similar to the previous example for CUSUM. We have again the incoming stream of observations and our reference sample, with a change at $t=200$ shown as the red dashed line. The KCUSUM algorithm is based on the linear statistic $P_{L}$, so at each time $n$, the MMD estimate $k(x_{n-1},x_{n})+k(y_{n-1},y_{n})-k(x_{n},y_{n-1})-k(x_{n-1},y_{n})$ is computed. Similar to the Cumulative Sum algorithm, a user defined threshold $h$ decides whether or not a change exists. With the stated parameters, a threshold of $h=5$ with drift $\delta=\frac{1}{40}$, found a change at $T_{KCUSUM}=225$.
@@
Let $T_{KCUSUM}$ be the change detector as described in Algorithm V.I. Let $p_{1}$ be the pre-change sample distribution. If we define $\delta>0$, the time to false alarm is at minimum:
@@
If $p_{2}$ is a distribution with $d_{k}(p_{1},p_{2})>\surd\delta$, then the worst case detection delay is at most:
@@
By substituting (14) into (16), we arrive at (11). To see the upper limits of ESADD, we take the sequence $\{x_{i}\}_{i\geq 1}$, which has a predetermined change on a non-even set, $t=2m-1$ for $m\geq 1$.
(17) $Z_{i}~{}\begin{cases}p_{1}\times p_{1}\times p_{1}\times p_{1}&\text{for $1\leq i<m$}\\ p_{2}\times p_{2}\times p_{1}\times p_{1}&\text{for $m\leq i$}\\ \end{cases}$
@@
An interesting case arises when a change occurs during an even sample. We consider a sequence that does not contain a rapid change, where samples are obtained from three distinct distributions:
(19) $Z_{i}~{}\begin{cases}p_{1}\times p_{1}\times p_{1}\times p_{1}&\text{for $1\leq i<m$}\\ p_{1}\times p_{2}\times p_{1}\times p_{1}&\text{for $i=m$}\\ p_{2}\times p_{2}\times p_{1}\times p_{1}&\text{for $m<i$}\end{cases}$
@@
By the assumption our kernel is bounded by $||k||_{inf}$, the specific bounds as described are maintained. Since the kernel is bounded, we also assume that $\delta$ is similarly bounded by $2||k||_{inf}$. Therefore, if $g_{\delta}(z)\leq 0$ for all $z$, then the algorithm cannot detect the change.
@@
Let us assume that Theorem V.II holds, and that $k$ and $g_{\delta}(z)$ are bounded by the previous proof. To determine ARL2FA,
@@
If $d_{k}(p_{1},p_{2})>\surd\delta$ describes the distribution $p_{2}$, then ARL2FA has an upper bound of:
(23) $ESADD_{KCUSUM}\leq\frac{2h}{d_{k}(p_{1},p_{2})^{2}-\delta}+\frac{8||k||^{2}_{\infty}}{(d_{k}(p_{1},p_{2})^{2}-\delta)^{2}}$
@@
to identify an $r$ to satisfy our conditions. To find the set of solutions, we expand our moment generating function $M(r)=E_{inf}[e^{rg_{\delta}(z)}]$: